OpenStudy (prince1342):

The tables below show the values of y corresponding to different values of x. Table A x 3 2 1 y 1 0 0 Table B x 3 5 5 y -2 1 2 Which statement is true for the tables? Both Table A and Table B represent functions. Both Table A and Table B do not represent functions. Table A does not represent a function but Table B represents a function. Table A represents a function but Table B does not represent a function.

OpenStudy (anonymous):

C

OpenStudy (prince1342):

For sure?

OpenStudy (anonymous):

yes it makes sense.let me know if u get it right

OpenStudy (prince1342):

thxx

OpenStudy (anonymous):

any time

OpenStudy (prince1342):

Is it alright if i post my other question here

OpenStudy (anonymous):

Yes

OpenStudy (prince1342):

Which of the following equations best represents the graph of f(p)? y = f(p) f(y) = p f(x) = p x = f(p)

OpenStudy (anonymous):

f(p) means that p is being inputted in to some function f. Therefore, if p is an input, the default expression for the output is y. Hence, inputting a value p into a function f would generate some output y. Therefore, the correct equation would be y = f(p). http://openstudy.com/study#/updates/526599a9e4b029b030da5084

OpenStudy (prince1342):

THANKSSSS

OpenStudy (prince1342):

If f(x) = 7x - 1, what does f(12) represent? The value of x when (7x - 1) = 12 The value of (7x - 1) when x = 12 The value of 12 f(x) when x = 12 The value of f(x) when (7x - 1) = 12

OpenStudy (prince1342):

If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5